OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (mertsj):

\[3\sqrt{2x}+x \sqrt{8x}-5\sqrt{18x}=3\sqrt{2x}+2x \sqrt{2x}-15\sqrt{2x}=-12\sqrt{2x}+2x \sqrt{2x}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (mertsj):

yw